OpenStudy (anonymous):

if f(x)=x^2+1 and g(x)=2x, find [f o g](x)

OpenStudy (anonymous):

f(g(x)) means that you must calculate g(x) and then put the result into f(x). For example, suppose f(x)=log(x) and g(x)=3x + 5. Then f(g(x))=log(3x+5). Does it sound more clear, now? You put g(x) everywhere you see x in the expression of f(x), to say it brutally

OpenStudy (anonymous):

so f(g(x))=2x+2^2?

OpenStudy (anonymous):

Mmmm, almost got it. g(x) = 2x, so you should put (2x) instead of x (in the expression of f(x)), with the parentheses, to be sure not to do distraction mistakes

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

would my answer be 2x^2+2?

OpenStudy (anonymous):

Remember not to forget the parentheses, they're fundamental! Also, the " +1 " in the expression of f(x) does not contain x, so you don't have to change it, it remains " + 1". Try again, now you will get it!

OpenStudy (anonymous):

2x^2+1

OpenStudy (anonymous):

Actually, the answer is 4 x^2 + 1, or (2x)^2 + 1. Those damn parentheses may change the result and mislead you! I hope to have been helpful :D (remember the parentheses!!!)

OpenStudy (anonymous):

thank you!